saleem bought 90 oranges at the rate of 3 oranges for rs 10 and sold them at the rate of 2 oranges for rs 9 find that saleem is in profit or loss and also find its percentage solution
Answer by Boreal(15235)

3 oranges for Rs10 is 90 oranges for Rs300, multiplying both by 30.
2 oranges for Rs9 is 90 oranges for Rs405, multiplying both by 45.
That is 105 Rs. Profit, or 35%, 105/300.
-------------
Check. He buys them for 10/3 and sells them for 9/2
9/2/10/3=27/20=1.35 or 35% increase.